OpenStudy (anonymous):

Show that the equations 8a-10b=1/3 and 20a-25b=5/6 Pls Help!!!

Directrix (directrix):

Show what about the equations? Perpendicular, parallel, or some other relationship?

OpenStudy (anonymous):

idk

Directrix (directrix):

Read the instructions that preceded the problem.

OpenStudy (anonymous):

oops I forgot to put have many solutions

Directrix (directrix):

*How* many solutions?

OpenStudy (anonymous):

It just says Show that the equations 8a-10b=1/3 and 20a-25b=5/6 have many solutions

Directrix (directrix):

These two equations are the same line. (8a - 10b = 1/3) -- multiply by 5/2 --> 20a -25b = 5/6 which is identical to the second equation. (20a - 25b = 5/6) In a sense, the two lines are one and intersect each other in every one of the infinitely many points on the line. That would be infinitely many solutions.

OpenStudy (anonymous):

Ok thanks :D